A potter's wheel with a moment of inertia of 8.0kg-m^2 has 5.0 N-m applied to it. It starts from rest. 10. seconds later it has gained what kinetic energy? (The answer is 78J, but I don't know how to arrive at the answer.)
Answer #1

T=I[angular accelaration]

= 5/8 = angular accelaration = 0.625

w = w0 + angluar accelaration * t

= w = 0.625*10

= 6.25v rad/sec

so the K.E = 0.5I w^2 = 0.5*8*6.25^2 = 156.625 J
Answer #2

Torque = I*alpha

alpha = 5/8 rad/s^2

w = w0 + alpha*t

w = 5/8*10 = 50/8 rad/s

Now,

KE = 1/2 *I*w^2 = 0.5*8*(50/8)^2 = 78 J
Answer #4

? = Qt/I = 5*10/8 = 6.25 rad/sec

2. KE =